# Refactoring Codebases through Library Design

### v1

@@ Abstract
Maintainable and general software allows developers to build robust applications efficiently, yet achieving these qualities often requires refactoring specialized solutions into reusable components. This challenge becomes particularly relevant as code agents become increasingly accurate at solving isolated programming problems. We investigate code agents’ capacity to refactor code in ways supporting growth and reusability. We present both a method and a benchmark for refactoring: Librarian, a sample-and-rerank method for generating reusable libraries, and MiniCode, a benchmark where code agents must minimize and refactor multiple independent solutions into a joint library. Compared to state-of-the-art code agents, Librarian achieves strong results on both compression and correctness on MiniCode, obtaining compression rates 1.6-2x better than coding agents while also improving correctness. We open-source our code and benchmark at https://code-refactor.github.io/.
@@ 1 Introduction
Writing code is mainly a matter of rewriting code: debugging, refactoring, optimizing, and other activities within the software engineering lifecycle. But poor rewrites incur technical debt, with such debt costing up to \$2 trillion annually [1]. This problem will likely worsen as language models become increasingly responsible for generating code, because they excel at solving isolated programming problems, but their context length demands a myopic view of the codebase. It is therefore valuable to understand not just the ability of language models to solve programming problems, but also their ability to rewrite and refactor code in ways that support growth and reuse.
Effective code refactoring at scale is a design problem. When refactoring codebases, developers must navigate design decisions around concerns such as generality, re-usability, and maintainability. A classic example illustrates this design challenge: Human programmers often create overly-specialized, redundant solutions to similar problems and would benefit from redesigning specialized solutions into a shared library. This consolidation requires careful design decisions about the right level of abstraction — neither too specific nor too general — and appropriate interfaces that balance flexibility with usability.
@@
Here we focus on refactoring multiple code sources into a reusable software library, and pose the following question: To what extent can code agents address this problem, both within human-written codebases, and also in language model-generated code? To answer that question, we develop a new method and a benchmark. This goes beyond past work [2, 3, 4, 5, 6, 7, 8] in library learning that synthesized subroutines across small programs in i.e. $\lambda$-calculus, instead tackling the more naturalistic problem of redesigning large bodies of code written in contemporary high-level languages, such as Python, producing classes, methods, and helper functions in the style of a human-written library. We propose a simple method, Librarian (Figure 1), which samples possible code rewrites then reranks those samples based on criteria designed to capture what it means to have a good refactoring. To generate potential rewrites, we develop methods for clustering pieces of code together that share common structure, so that a succinct prompt can rewrite them jointly into their refactored form.
To evaluate our method and systematically assess the capability of current systems to perform such design-intensive refactorings, we introduce a new benchmark, MiniCode. Existing benchmarks such as SWE-Bench [9], Commit0 [10], and RefactorBench [11] primarily focus on functional correctness and do not directly evaluate an agent’s ability to perform large-scale code consolidation or library design. MiniCode fills this gap by explicitly challenging code agents to minimize and refactor multiple specialized code sources into a unified library. MiniCode requires agents to design libraries that expose a general interface shared across multiple use cases in two domains: competition coding and repository-level Python applications. This involves optimizing not only for functional correctness but also for broader software engineering objectives such as reusability and maintainability, making the task substantially more open-ended than prior benchmarks.
Our results show that state-of-the-art code agents, based on Claude 3.7 Sonnet and o4-mini, struggle to jointly preserve correctness and improve reusability across both domains of MiniCode. In the competition coding domain, our method Librarian improves refactoring quality by 1.89x while also enhancing correctness. However, on the repository-level refactoring, even the strongest agents fail to produce high-quality refactorings, highlighting a substantial gap between current capabilities and the demands of design-oriented code rewriting. Addressing this challenge remains an open and important direction for future research.
Figure 1: Overview of the problem that we study and the general structure of its solutions. Given a collection of different code sources, where a source is either program or repository, we design a modular and reusable library. To do this we cluster together related programs into tuples (left), sample different rewrites using a language model (right), and select the rewrites which optimize various criteria, such as compression, while validating candidate rewrites using test cases (pass rate).
@@ 2 Related work
Repo-level language model coding agents.
Recent work has explored the application of language models to repository-level software engineering tasks. Existing benchmarks include SWE-bench [9], which evaluates models on their ability to resolve real-life GitHub issues, and Commit-0 [10], which challenges AI systems to reconstruct entire code repositories from scratch. Such benchmarks primarily evaluate functional correctness via unit tests, without assessing the quality or maintainability of the resulting codebase. RefactorBench [11] takes a step in this direction by benchmarking the ability to follow specific refactoring instructions. Our work differs in that we expect models to perform a more open-ended task: Redesigning code to be more modular and compact by discovering and drawing out reused abstractions.
@@
Systems which perform library learning research discover shared abstractions across a large number of small programs, which they use to automatically define new subroutines. Systems such as DreamCoder [4], Trove [12], LiLo [7], and REGAL [3] automatically construct such libraries with the goal of making future program synthesis tasks easier to solve, once the learned library is in hand. Our work is closest to REGAL [3], which clusters related code and refactors using language models. However, existing library learning approaches have primarily been demonstrated in small-scale, constrained domains, limiting their applicability to typical software engineering tasks, such as consolidating multiple repositories into cohesive libraries. By framing library learning within the context of realistic, large-scale code repository development, we expand the relevance of library learning to everyday software engineering practice.
@@
While our goal is to optimize the quality of libraries, other works focus on improving code speed through correctness-preserving transformations [13, 14, 15]. Both forms of program optimization, compression and speed, require more exploration than only optimizing for correctness, as there does not exist a ground-truth answer. Prior work on program optimization benchmarks study code at the file level. We propose a benchmark that transforms programs at a larger scale, across multiple code repositories.
@@ 3 Problem Statement
In this section, we propose a refactoring task: Given multiple code sources that contain problem-specific implementations, the goal is to create a cohesive library that captures shared abstractions. This library must reduce the total code size while supporting all original use cases, potentially opening up new use cases as well by mining and formalizing latent shared abstractions. This is accomplished by searching for refactorings that are both correct and simple. Correctness is straightforward to define as the fraction of unit tests passed, but simplicity is more elusive.
One potential measure of simplicity is counting the total number of tokens in the proposed library and refactored code. However, just minimizing program size has obvious failure modes: code should also be natural, elegant, and extensible, which can be in tension with merely finding the shortest program. [^1] Perl Golf is a game where participants attempt to write the shortest Perl program accomplishing a given task. The resulting code is famously incomprehensible, even by the standards of Perl. To address these concerns, we follow prior work in program synthesis [16, 6, 17, 18] and define simplicity as the minimum description length (MDL), or negative log probability under a reference distribution.
Formally, we are given a collection of code sources $\left\{\rho_{n}\right\}_{n=1}^{N}$, and output both a new library $\mathcal{L}$, as well as rewritten refactorings of the original code sources, $\left\{\rho^{\prime}_{n}\right\}_{n=1}^{N}$. We define the pass rate $\tau(\rho_{n})$ as the fraction of unit tests program $\rho_{n}$ passes. In practice we are concerned both with the case where we are refactoring several code sources ($N>1$) and also the case where there is only a single large code source we are refactoring ($N=1$).
We optimize the following objective, which rewards refactorings that pass at least as many tests as the original program and minimize MDL:
@@
(,$\ell\left(\mathcal{L},\left\{\rho_{n}^{\prime}\right\}\right)=\begin{cases}-\log p_{\text{LM}}(\mathcal{L})+\sum_{n}-\log p_{\text{LM}}(\rho_{n}^{\prime}\mid\mathcal{L})&\forall\rho_{n},\tau(\rho_{n})\leq\tau(\rho_{n}^{\prime})\\ \infty&\text{otherwise}\end{cases}$ (1)
where $p_{\text{LM}}(\rho_{n}^{\prime}|\mathcal{L})$ is the probability of the suffix $\rho_{n}^{\prime}$ given the prefix $\mathcal{L}$ under a language model, effectively concatenating the library and the program into one prompt, but only counting the perplexity of the later program tokens.
@@ 4 Librarian: Refactoring Code to Create Libraries
This section details our method to compress collections of code sources into libraries, while migrating the code sources to use these shared building blocks. Figure 1 illustrates our method, Librarian.
Librarian follows a simple sample-and-rerank framework to maximize our refactoring objective described in Section 3. It maintains and grows a library of useful functions as part of this objective.
Concretely, our framework follows:

### v3

@@ Abstract
Maintainable and general software allows developers to build robust applications efficiently, yet achieving these qualities often requires refactoring specialized solutions into reusable components. This challenge becomes particularly relevant as code agents become used to solve isolated one-off programming problems. We investigate code agents’ capacity to refactor code in ways that support growth and reusability. We first investigate what makes a good refactoring, finding via simulation results and a human study that Minimum Description Length best correlates with preferable refactorings. We then present both a benchmark and a method for refactoring: MiniCode, a benchmark where multiple files must be refactored into a shared library, and Librarian, a sample-and-rerank method for generating reusable libraries. We compare Librarian to state-of-the-art library generation methods, and study it on real-world code bases.
@@ 1 Introduction
Writing code is mainly a matter of rewriting code: debugging, refactoring, optimizing, and other activities within the software engineering lifecycle. But poor rewrites incur technical debt, with such debt costing up to \$2 trillion annually (Tews, ). This problem will likely worsen as language models become increasingly responsible for generating code, because they excel at solving isolated programming problems, but their context length demands a myopic view of the codebase. Effective code refactoring at scale is a design problem whose concerns center on re-usability and maintainability. A classic example illustrates this design challenge: Human programmers often create overly-specialized, redundant solutions to similar problems and would benefit from redesigning specialized solutions into a shared library.
@@
Here we focus on refactoring multiple files into a reusable software library, which raises two questions: (1) How should we quantify the quality of a library refactoring, and (2) How can we use LLMs to refactor specialized code into reusable libraries? To answer those question, we develop a new method and a benchmark. This goes beyond past work in library learning (Wong et al., 2021; Stengel-Eskin et al., 2024; Ellis et al., 2021; Bowers et al., 2023; Dechter et al., 2013; Grand et al., 2023; Liang et al., 2010b) which synthesizes subroutines across small programs in i.e. $\lambda$-calculus, because instead we address the more naturalistic problem of redesigning code written in contemporary high-level languages, such as Python, producing classes, methods, and helper functions in the style of a human-written library. We develop a method, Librarian (Figure˜1), which samples possible code rewrites and then reranks those samples based on criteria designed to capture a good refactoring. To generate potential rewrites, we develop methods for clustering pieces of code together that share common structure so that a succinct prompt can rewrite them jointly into their refactored form. To find strong criteria for ranking potential rewrites, we study a variety of metrics across machine learning and software engineering, both on programming benchmarks and via a human study.
To evaluate our method and systematically assess the capability of current agents to generate libraries, we introduce a new benchmark, MiniCode, which addresses three key desiderata missing from existing benchmarks. First, open-ended design: unlike SWE-Bench (Jimenez et al., 2024), Commit0 (Zhao et al., 2025), and RefactorBench (Gautam et al., 2025) which primarily focus on functional correctness, MiniCode presents an unconstrained library design problem. Agents create a library that can be imported back into a repository, with complete freedom to design the interface and implementation from scratch—optimizing for software engineering objectives like reusability and maintainability. Second, verifiability: we ensure objective evaluation by retaining the unit tests from all repositories that will import the designed library, allowing us to verify that the solutions work correctly across multiple use cases. Third, large context: agents must understand and synthesize information from multiple code sources (files) simultaneously to design a unified library that consolidates specialized code sources into a general interface.
Figure 1: Overview of the refactoring problem. A refactoring task comprises a set of files. We refactor the files by designing a new library. Candidate refactorings are evaluated based on a refactoring metric, and are expected to maintain correctness of the original code sources (pass rate). We explore several refactoring metrics in this paper.
We contribute the following:
  1. 1.
    Study of different metrics for library creation via a user study and program synthesis benchmarks, with the surprising finding that neural language models offer a stronger signal of refactoring quality than classic metrics from the software engineering community.
  2. 2.
    A new benchmark, MiniCode, covering both competition programming and real-world machine learning programs from the Transformers and Diffusers libraries.
  3. 3.
    A new algorithm, Librarian, which outperforms prior library learners on existing benchmarks and scales to complex codebases.
  4. 4.
    Refactorings of real-world codebases widely used in the machine learning community. Our method successfully refactors the Transformers and Diffusers libraries from Huggingface to be shorter and more reusable. To the best of our knowledge this is the first time library learning has have been successfully applied to real-world software projects.
@@ 2 Related work
Systems which perform library learning research discover shared abstractions across a large number of small programs, which they use to automatically define new subroutines. Systems such as DreamCoder (Ellis et al., 2021), Trove (Wang et al., 2024), LiLo (Grand et al., 2023), and REGAL (Stengel-Eskin et al., 2024) automatically construct such libraries with the goal of making future program synthesis tasks easier to solve, once the learned library is in hand. Our work is closest to REGAL (Stengel-Eskin et al., 2024), which clusters related code and refactors using language models. However, existing library learning approaches have primarily been demonstrated in small-scale, constrained domains, limiting their applicability to typical software engineering tasks, such as consolidating multiple repositories into cohesive libraries. By framing library learning within the context of realistic software development, we expand the relevance of library learning to the everyday practice of software engineering.
Repo-level coding benchmarks.
Recent work has explored the application of language models to repository-level software engineering tasks. Existing benchmarks include SWE-bench (Jimenez et al., 2024), which evaluates models on their ability to resolve real-life GitHub issues, and Commit-0 (Zhao et al., 2025), which requires agents to fill in function definitions. Such benchmarks primarily evaluate functional correctness via unit tests, without assessing the quality or maintainability of the resulting codebase. RefactorBench (Gautam et al., 2025) takes a step in this direction by benchmarking the ability to follow specific refactoring instructions. Our work differs by requiring models to perform a more open-ended task: Redesigning code to be more modular and compact by discovering reused abstractions, while retaining verifiability by re-using downstream unit tests. Additionally, libraries must be created without any scaffolding limitations such as preexisting function definitions affording more design freedom than Commit-0.
@@
While our goal is to optimize the quality of libraries, other works focus on improving execution speed through correctness-preserving transformations (Waghjale et al., 2024; Ouyang et al., 2025; Schkufza et al., 2013). Both forms of program optimization, compression and speed, are more open-ended than optimizing only for correctness, as there does not exist a ground-truth answer. Prior work on program optimization benchmarks study code at the file level. We propose a benchmark that transforms programs at a larger scale, across multiple code files.
@@ 3 Problem Statement
Given $N$ related files $\left\{\rho_{n}\right\}_{n=1}^{N}$, the goal is to create a library $\mathcal{L}$ that captures shared abstractions. The original files $\left\{\rho_{n}\right\}_{n=1}^{N}$, which define the refactoring problem, we call a task. The new library must support all original use cases in the task by extracting latent shared abstractions. This is accomplished by searching for refactorings that are both correct and ‘natural’. Correctness is straightforward to define via unit tests, but naturalness is more challenging to quantify.
Shorter code is potentially simpler and less redundant. One potential metric is to count the number of tokens, lines-of-code, or syntax tree nodes in the proposed library and refactored code  (Dechter et al., 2013; Polozov & Gulwani, 2015; Bowers et al., 2023; Cao et al., 2023). But minimizing program size has obvious failure modes: code should also be understandable and extensible, which can be in tension with merely finding the shortest program. [^1] Perl Golf is a game where participants attempt to write the shortest Perl program accomplishing a given task. The resulting code is famously incomprehensible, even by the standards of Perl. Other work in program synthesis (Liang et al., 2010a; Solomonoff, 1964; Ellis et al., 2021) instead optimizes Minimum Description Length (MDL), or negative log probability under a reference distribution. In the software engineering community, other metrics such as cyclomatic complexity and maintainability index have been defined for similar purposes: These are more complex metrics that examine the syntax tree, call graph, and other statically-analyzable structures (McCabe, 1976). What metric should we use? We revisit this question in Section 6, where we empirically compare candidate metrics and human preferences before fixing our choice for the rest of the paper.
For now assume a placeholder metric $M$ measuring refactoring quality; we seek to minimize $M$ while preserving correctness. Given a task comprising files $\left\{\rho_{n}\right\}_{n=1}^{N}$, we output both a new library $\mathcal{L}$, as well as rewritten refactorings of the original files, $\left\{\rho^{\prime}_{n}\right\}_{n=1}^{N}$. We define tests passed $\tau(\rho_{n})$ as the set of unit tests $\rho_{n}$ passes, and consider both refactoring several files ($N>1$) and also refactoring a single large file ($N=1$). We optimize the following objective, which prefers passing at least the same tests as the original programs and minimizing the chosen metric $M$:
@@
$\ell\left(\mathcal{L},\{\rho_{n}^{\prime}\}\right)=\begin{cases}M(\mathcal{L},\{\rho_{n}^{\prime}\})&\forall\rho_{n},\tau(\rho_{n})\subseteq\tau(\rho_{n}^{\prime})\\ \infty&\text{otherwise}\end{cases}$ (1)
4 MiniCode—Library Design and Refactoring Benchmark
MiniCode presents systems with a task comprising a set of files, then asks them to refactor the files into a unified library alongside refactorings of the original files. There are two key desiderata for benchmark tasks: They should have related files sharing latent abstractions, and should also be verifiable, to measure how well refactored files preserve functional correctness. We source a variety of problems (Table˜1).
Table 1: MiniCode Statistics
Domain Files Tasks Avg LoC Avg Tests / file
Code Contests (Li et al., 2022b) 300 10 87 10
Transformers (Wolf et al., 2020) 10 1 538 181
Diffusers (von Platen et al., 2022) 11 2 685 75
Logo (Wong et al., 2021) 300 1 10 1
Date (Srivastava et al., 2023) 246 1 14 1
CodeContests.
Competition problems are crafted with specific variations of algorithmic approaches in mind, resulting in both shared latent concepts and the required test cases. As a result, competition coding is both verifiable, and ready to refactor. We therefore take solutions, prompts, and tests from CodeContests (Li et al., 2022a), a competition programming dataset.
Huggingface Transformers Library.
We test refactoring across implementations of large language and vision–language models from the Huggingface transformers repository (modelling_<name>.py files, e.g., Qwen2, LLaMA, DeepSeek-V3). Unlike competition coding, these sources are production-scale and Huggingface requires that all changes pass an extensive suite of integration tests before merging into the main branch. A refactoring is only deemed correct if it passes the unmodified Transformers test suite, making this a high-stakes setting that requires correctness and compatibility.
Huggingface Diffusers Library.
We test refactoring across implementations of diffusion models from the Huggingface diffusers repository (unet_<name>.py and scheduler_<name>.py files, e.g., Stable Diffusion UNet, DDPMScheduler), yielding two distinct tasks. Like Transformers, Diffusers requires that all changes pass a comprehensive suite of integration tests before merging into the main branch.
Logo & Date.
The library learning literature already has existing benchmarks: Typically they seek to learn a single library from a task comprising many sources, and then test that library on holdout program synthesis tasks. Logo and Date were used in the recent related work REGAL (Stengel-Eskin et al., 2024), which we incorporate wholesale to understand how our new method compares to state-of-the-art library learning. The associated programming problems were created by humans, but their solutions were generated by gpt-3.5-turbo.
@@ 5 Librarian: Refactoring Code to Create Libraries
Librarian generates a new library from a set of files, while migrating the files to use that new library (Figure 1), following a sample-and-rerank framework: Prompting a backend LLM or agent to sample $K$ candidates, and picking the one minimizing the loss $\ell$. Naively,
$\mathcal{L},\left\{\rho^{\prime}_{n}\right\}=\operatorname*{arg\,min}_{\mathcal{L},\left\{\rho^{\prime}_{n}\right\}\in\textsc{Sample}_{K}\left(\left\{\rho_{n}\right\}\right)}\ell\left(\mathcal{L},\left\{\rho^{\prime}_{n}\right\}\right)$ (2)
for metric $M$ and sampling budget $K$. But this cannot work for large tasks with many programs, which would not fit into the context of most LLMs. Even long context models cannot process the entirety of e.g the Linux kernel, and even if they could, it is not clear that such a strategy is the most efficient way of focusing the language model’s attention. To address this, we wrap sample-and-rerank with a clustering algorithm that decomposes the task into manageable chunks, described next.
Clustering.
Meaningful abstractions arise when programs share underlying functionality or structure. To surface these, we cluster the task’s files into small groups that are likely to share reusable structure, and refactor each cluster separately from the rest. This decomposition shrinks the prompt size, and gives independent searches for the best per-cluster refactoring, which may be more tractable.
Librarian’s clustering extends REGAL (Stengel-Eskin et al., 2024), which clusters programs by assuming each program is paired with a natural language description of the problem it solves, and clustering embeddings of those descriptions. Since similar problems need not imply similar solution code, we instead prompt a model to summarize each file and cluster by these summaries. Specifically, we define $\textsc{Cluster}_{S}\left(\left\{\rho_{n}\right\}\right)$ as performing agglomerative clustering (Ward Jr, 1963) on the task’s files $\left\{\rho_{n}\right\}$ to produce a set of set of files, each of which is a cluster of size $S$. We use text-embedding-ada-002 to embed descriptions of code sources for clustering.
Combining clustering with sample-and-rerank.
The simplest approach is to refactor each cluster independently and take the union (concatenation) of each cluster’s library:
$\mathcal{L}^{\star}$ $=\bigcup_{c\in\textsc{Cluster}_{S}(\left\{\rho_{n}\right\})}\mathcal{L}_{c}$ (3)
$\left\{\rho^{\prime}_{n}\right\}$ $=\bigcup_{c\in\textsc{Cluster}_{S}(\left\{\rho_{n}\right\})}\left\{\rho^{\prime}_{c,i}\right\}_{i=1}^{|c|}$ (4)